I'll rate as soon as possible, thanks!
A 7.6 cm high flower placed 1.00 m from +50.0 mm l
Answer #1

1/f = 1//do + 1/d1

1/50 = 1/1000 + 1/d1 [ here do is the distance of object)

d1 = 52.63 mm

it is positive so image is real

M = - di/do =-50/52.63 = -0.95

height of image = -0.95 x 7.6 cm = -7.22 cm

So image is inverted since negative sign.
Answer #2

1/f = 1//do + 1/d1

1/50 = 1/1000 + 1/d1 [ here do is the distance of object)

d1 = 52.63 mm

it is positive so image is real

hi/ho = M = -di/do => hi/76 = -52.63/1000 = hi = -(76 * 52.63)/1000

hi = -4mm

So image is inverted since negative sign.

answered by: Interested